given that y1(x) = x^-1 is one solution of the differential equation 2x^2 dy^2/dx^2 + 3xdy/dx - y=0, x>0 find a second linearly independent solution of the equation
1
Expert's answer
2020-04-10T17:21:57-0400

Look for a solution in the form "y=y_1z=\frac{z}{x}", then "y'=\frac{z'}{x}-\frac{z}{x^2}" and "y''=z''\left(\frac{1}{x}\right)+2z'\left(\frac{1}{x}\right)'+z\left(\frac{1}{x}\right)''=\frac{z''}{x}-\frac{2z'}{x^2}+\frac{2z}{x^3}"

Then "0=2x^2\left(\frac{z''}{x}-\frac{2z'}{x^2}+\frac{2z}{x^3}\right)+3x\left(\frac{z'}{x}-\frac{z}{x^2}\right)-\frac{z}{x}=2z''x-z'"

Let "z'=u", then "2u'x-u=0" or "\frac{2u'ux-u^2x'}{x^2}=0".

Since "2u'u=(u^2)'", we have "0=\frac{2u'ux-u^2x'}{x^2}=\frac{(u^2)'x-u^2x'}{x^2}=\left(\frac{u^2}{x}\right)'", so "\frac{u^2}{x}=C"

Then "u=\sqrt{Cx}=\sqrt{|C|}\sqrt{x}" if "C\ge 0", or "u=\sqrt{Cx}=\sqrt{|C|}\sqrt{-x}" if "C\le 0".

Denote "\sqrt{|C|}" by "D", then "u=D\sqrt{x}" or "u=D\sqrt{-x}"

Replace "u" by "z'": "z'=D\sqrt{x}" or "z'=D\sqrt{-x}".

Then "z=\frac{2}{3}D\sqrt{x^3}+A" or "z=-\frac{2}{3}D\sqrt{-x^3}+A". In every case "z=B\sqrt{|x|^3}+A".

Since "y=\frac{z}{x}", we have "y=Bs(x)\sqrt{|x|}+\frac{A}{x}=E\sqrt{|x|}+\frac{A}{x}", where "s(x)=\begin{cases}
1,&\text{if $x>0$}\\
0,&\text{if $x=0$}\\
-1,&\text{if $x<0$}
\end{cases}"

So "\sqrt{|x|}" is the second solution.

Answer: "\sqrt{|x|}"